### v2

@@ Appendix A Related Work
Visual instruction tuning is a key technique for enhancing multimodal large language models by aligning visual inputs with textual instructions to improve understanding and generation tasks Liu et al. (2023b). Traditionally, these instructions are built using English-language data from visual question answering and other datasets Liu et al. (2023b; a); Xu et al. (2023); Liu et al. (2024); Tong et al. (2024); Beyer et al. (2024); Zhu et al. (2024); Dai et al. (2023); Tong et al. (2024); Li et al. (2024d). Researchers often supplement this with synthetic instruction tuning data, generating large volumes of instructional pairs to possibly cover multiple languages too Geigle et al. (2024b); Li et al. (2023; 2024d). However, these instruction-tuning datasets have mostly been task-focused and lack conversational capabilities. Further, while translation gives lends to multilingual capabilities, the data remains to be culturally homogeneous. By curating multilingual and multicultural instruction tuning data across various task types, our model is designed to intuitively understand and engage with users from diverse demographics.
@@
Multilingual MLLMs have evolved from dual-encoder-based models, only capable of understanding and reasoning Ni et al. (2021); Zeng et al. (2022); Jain et al. (2021), to encoder-decoder models capable of multilingual text generation as well Shan et al. (2022); Chen et al. (2023); Geigle et al. (2024b). Despite their advancements, these models have remained focused on conventional tasks such as VQA and image captioning. Moreover, most efforts have centered around training with multilingual text, while little attention has been given to curating culturally diverse image datasets. Even for text, despite the focus on multilinguality, few attempts have been made to reflect cultural diversity in instructions and captions. As a result, these models tend to reflect a Western-centric bias. By selecting culturally diverse images from LAION and intentionally integrating this diversity into our instructions and captions, our model aims to serve a wide range of users in an inclusive and equitable manner.
@@ Appendix F Qualitative Examples from xChatBench
We show some interesting qualitative examples from xChatBench, including good cases in Figure 9, Figure 10, Figure 11, Figure 12, Figure 13, Figure 14, as well as bad cases in Figure 15, Figure 16. The examples included the scoring rubric, query, response from our Pangea, reference answer, and LLM-as-Judge feedback.

### v3

@@ 6 Conclusion
Acknowledgments
This work was supported in part by a Carnegie Bosch Institute Fellowship to Xiang Yue, as well as grants from DSTA Singapore, and the Programs for Bridging the gap between R&D and the IDeal society (society 5.0) and Generating Economic and social value (BRIDGE)/Practical Global Research in the AI × Robotics Services, implemented by the Cabinet Office, Government of Japan. The training is supported by the CMU FLAME Center. The authors would like to thank Google Gemini credits for data construction and evaluation. The authors would also like to thank CMU NeuLab colleagues for their constructive comments.
@@ Appendix A Related Work
Visual instruction tuning is a key technique for enhancing multimodal large language models by aligning visual inputs with textual instructions to improve understanding and generation tasks (Liu et al., 2023b). Traditionally, these instructions are built using English-language data from visual question answering and other datasets (Liu et al., 2023b; a; Xu et al., 2023; Liu et al., 2024; Tong et al., 2024; Beyer et al., 2024; Zhu et al., 2024; Dai et al., 2023; Tong et al., 2024; Li et al., 2024d). Researchers often supplement this with synthetic instruction tuning data, generating large volumes of instructional pairs to possibly cover multiple languages too (Geigle et al., 2024b; Li et al., 2023; 2024d). However, these instruction-tuning datasets have mostly been task-focused and lack conversational capabilities. Further, while translation gives lends to multilingual capabilities, the data remains to be culturally homogeneous. By curating multilingual and multicultural instruction tuning data across various task types, our model is designed to intuitively understand and engage with users from diverse demographics.
@@
Multilingual MLLMs have evolved from dual-encoder-based models, only capable of understanding and reasoning (Ni et al., 2021; Zeng et al., 2022; Jain et al., 2021), to encoder-decoder models capable of multilingual text generation as well (Shan et al., 2022; Chen et al., 2023; Geigle et al., 2024b). Despite their advancements, these models have remained focused on conventional tasks such as VQA and image captioning. Moreover, most efforts have centered around training with multilingual text, while little attention has been given to curating culturally diverse image datasets. Even for text, despite the focus on multilinguality, few attempts have been made to reflect cultural diversity in instructions and captions. As a result, these models tend to reflect a Western-centric bias. By selecting culturally diverse images from LAION and intentionally integrating this diversity into our instructions and captions, our model aims to serve a wide range of users in an inclusive and equitable manner.
@@ Appendix F Qualitative Examples from xChatBench
One important application of MLLMs is to answer users’ queries in the wild. Here, we show the outputs of Pangea for the multimodal chat queries from our xChatBench. The examples included the scoring rubric, query, response from our Pangea, reference answer, and LLM-as-Judge feedback. As shown in Appendix Figure 9, 10, 11, 12, 13, 14, Pangea successfully interprets the figures in different tasks and generates fluent and readable in certain languages. These qualitative examples further demonstrate the remarkable visual understanding ability of Pangea in multilingual contexts. On the other hand, we also identified a few bad cases shown in Figure 15, 16. Despite generating relevant responses to the queries, the model does not capture the key details of the images due to the lack of knowledge, which points out potential improvement directions in the future.